Using the length of the day, compute the speed of a object on the earth's equator due to the earth's rotation (=6,370km)


1
Expert's answer
2021-04-29T10:46:21-0400

The speed is distance over time. The distance is the length of the equator circle:


"d=2\pi R,\\\space\\
v=\frac dt=\frac{2\pi R}{t},\\\space\\
v=\frac{2·3.14·6370}{24}=1668\text{ km/h}."